An express train travels 150 miles in the same time that a freight train travels
100 miles. If the express goes 20 miles per hour faster than the freight, find the
rate each train travels

Answer by mananth(16949)

An express train travels 150 miles in the same time that a freight train travels
100 miles. If the express goes 20 miles per hour faster than the freight, find the rate each train travels
..
freight train - speed = x , d=100 miles
express train speed = (x+20),d= 150miles
times are same
100/x= 150/(x+20)
100(x+20)=150x
100x+2000=150x
150x-100x=2000
50x = 2000
x= 40 mph freight train speed
express train speed = x+20 = 60 mph
...
m.ananth@hotmail.ca